# HDCompression: Hybrid-Diffusion Image Compression for Ultra-Low Bitrates thanks: Equal Contribution

### v2

@@ III-B1 LIC-assisted indices map prediction
where $enc_{\mathbf{v}}$, $enc_{id_{i}}$, and $enc_{mask}$ correspond to the encoded DRV, encoded $id_{i}$, and encoded $mask$ respectively. To fully utilize the encoded global information from DRV, we replace all $enc_{mask}$ with $enc_{\mathbf{v}}$ in $\textbf{enc}_{P}$ before feeding it to the transformer decoder, so that all masked locations have global priors for better token prediction.
Figure 4: VQ Correction Module for dual-stream merging.
Fig. 3 illustrates the detailed process of DRV $\textbf{v}_{joint_{P}}$-guided masked token prediction using T. Note that the 2D VQ indices map $\textbf{d}\in\mathbb{R}^{\frac{H}{m}\times\frac{W}{m}}$ is first flattened into a sequence before being fed into T in transformer. Since LIC-assisted Indices Map Prediction is a form of conditional generation and high-level global priors are already fused into the masked locations via the $\textbf{v}{joint_{P}}$ embedding, we find it beneficial to further constrain the prediction output to more closely align with the ground truth. This is achieved by integrating lower-level fidelity features from the enhanced LIC output $\hat{\textbf{x}}_{lic}$ into the transformer decoder. Specifically, inspired by the “Prediction Assistance” design in HybridFlow [22], we use a forked VQ Encoder $\textbf{E}_{P}$, with the output channel $\textbf{C}_{vq}$ modified to match the hidden dimension $h_{dim}$ of the transformer, to extract a lower-level feature map $\textbf{fm}_{P}\in\mathbb{R}^{\frac{H}{n}\times\frac{W}{n}\times h_{dim}}$ from $\hat{\textbf{x}}_{lic}$. The feature map $\textbf{fm}_{P}$ is then flattened to the shape $\frac{HW}{n^{2}}\times h_{dim}$ and fed into the cross-attention modules of the transformer decoder, serving as the $key$ and $value$.
To avoid transmitting $\textbf{v}_{joint_{P}}$ and save bits, we use a diffusion process, similar to the approach of using ground-truth infused DRV $\textbf{v}_{gt}$ for enhancing the LIC stream. We sample a $\hat{\textbf{v}}_{joint_{P}}$ in decoder from a noisy initialization, conditioned on a vector $\textbf{c}_{P}$ extracted from $\hat{\textbf{x}}_{lic}$ by following Eq.(2$\sim$4). The denoising module for generating $\hat{\textbf{v}}_{joint_{P}}$ has the same simple network structure and sampling scheduler as the denoising module for generating $\hat{\textbf{v}}_{joint_{E}}$, and the network for extracting $\textbf{c}_{P}$ from $\hat{\textbf{x}}$ share the same architecture with network $E_{C}$ for extracting c from $\hat{\textbf{x}}$ in the LIC stream.
III-B2 LIC-assisted conditioned pixel decoding
@@
It is worth mentioning that our method has the identical bit consumption as the dual-stream HybridFlow [22], since both the ground-truth DRV $\textbf{v}_{gt}$ for LIC stream enhancement and latent $\textbf{v}_{join_{P}}$ for improved indices prediction are reconstructed in the decoder. In other words, by learning diffusion priors, we enhance dual-stream performance without additional transmission overhead.
III-C Training Pipeline
@@
III-C1 Basic flow pre-training
The LIC stream uses the pre-trained LIC encoder from MLIC. The VQ-Encoder $\textbf{E}_{vq}$ and the learned visual codebook C use the pre-trained VQGAN model [17]. These pre-trained components are designed to either pursue high-fidelity reconstruction or high-quality reconstruction, ensuring the baseline performance of our dual-stream system.
III-C2 DRV-based enhancement module training
@@
III-C3 DRV-based transformer predictor training
@@
III-C4 VQ correction module training
@@
III-C5 DRV-diffusion module training
@@
IV Experiments
Datasets. Our HDCompression model is trained on ImageNet-1k [37], with 1000 categories. In alignment with HybridFlow [22], performance is evaluated over three benchmarks: Kodak [38], CLIC 2020 test set [39] and Tecnick dataset [40].
Model configurations. Training images are cropped into $256\!\times\!256$ patches. To achieve approximately 0.025 bpp using pre-trained MLIC models [6] as the Base LIC, which typically provide a minimum of 0.1 bpp, input patches for the LIC stream are further downsized to $128\times 128$ via bilinear interpolation. Both streams have the same downsampling factor $m=n=16$. The lightweight diffusion modules use a 4-step DDPM scheduler. We unify the loss weights as follows: $w_{1}=1.2$ for $L_{1}$ pixel loss, $w_{2}=0.8$ for AlexNet-based LPIPS perceptual loss, and $w_{3}=0.12$ for UNet-based pixel-wise discriminator GAN loss.
Compared baselines. We compare with several representative image compression methods: 1) Traditional hand-crafted VVC & BPG; 2) MLIC [6], a single-streamed conventional LIC method; 3) VQGAN [17] & Fine-tuned VQGAN [41], single-streamed VQ-codebook-based methods; and 4) HybridFlow [22], a dual-stream framework that straightforwardly combines pre-trained LIC with VQ-codebook-based stream. To obtain ultra-low bitrates, we set QP ranging in $[45,51]$ for VVC & BPG, and downsize the input image (1/2 width and height) for MLIC.
Evaluation metrics. We evaluate commonly used PSNR and LPIPS [42]. PSNR measures pixel-level distortion and LPIPS assesses the visual quality.
IV-A Quantitative Results
HDCompression has the same bpp as HybridFlow, which operates over $[0.025,0.065]$ bpp range, spanning from fully masked indices map (lowest quality) to unmasked indices map (highest quality). As shown in Fig. 5, traditional handcrafted VVC & BPG and conventional MLIC outperform codebook-based methods over PSNR, due to their learning target of minimizing pixel-level distortions. However, codebook-based methods perform significantly better for perceptual LPIPS. The distortion-driven focus gives artificially inflated PSNR, which omits image details and prefers overly smoothed regions. In contrast, single-stream codebook-based methods emphasize LPIPS, neglecting fidelity to the original image, resulting in $>$4 dB PSNR drop compared with traditional methods at the same bpp.
@@ IV-A Quantitative Results
Figure 6: Qualitative Comparison of our method to the baselines. ”1_4” mask strategy ($75\%$ mask ratio on $\textbf{d}_{vq}$) is utilized to maintain around 0.035 bpp within the similar range of the compared baselines. Zoom in for better visualization.
IV-B Qualitative results
@@
IV-C Ablation Study on Hybrid-Diffusion Modules
@@
IV-C1 DRV-based enhancement for the LIC stream
We compare our LIC stream output ${\hat{\textbf{x}}}_{lic}$ of incorporating the enhancement module against the original output ${\hat{\textbf{x}}}$ of the pre-trained MLIC. As illustrated in Fig. 7 (a), the enhancement module improves the quality of the original LIC output ${\hat{\textbf{x}}}$, particularly by reducing blurs. The enhancement module significantly enhances LPIPS of ${\hat{\textbf{x}}}$ while maintaining almost the same PSNR across various datasets in Fig. 7 (b).
IV-C2 DRV-based transformer for mask prediction
We compare the logit-wise token prediction loss (Eq. 10) of the DRV-based transformer mask predictor against the naive transformer mask predictor in HybridFlow without using DRV. As shown in Fig. 7 (b) where the indices map is masked by ”1_4” masking schedule (75% masking ratio), the prediction loss is dropped by 18.5% on average by using DRV in the transformer encoder, leading to 15% accuracy improvement in token prediction on average. Thus, more ground-truth indices are recovered to provide more specific details to the generative stream that might be neglected by the LIC stream.
IV-C3 VQ correction for dual-stream merging

### v3

@@ 3.2.1 LIC-assisted indices map prediction
where $enc_{\mathbf{v}}$, $enc_{id_{i}}$, and $enc_{mask}$ correspond to the encoded DRV, encoded $id_{i}$, and encoded maskmask respectively. To fully utilize the encoded global information from DRV, we replace all $enc_{mask}$ with $enc_{\mathbf{v}}$ in $\textbf{enc}_{P}$ before feeding it to the transformer decoder, so that all masked locations have global priors for better token prediction.
Fig. 3 illustrates the detailed process of DRV $\textbf{v}_{joint_{P}}$-guided masked token prediction using T. Note that the 2D VQ indices map $\textbf{d}\in\mathbb{R}^{\frac{H}{m}\times\frac{W}{m}}$ is first flattened into a sequence before being fed into T in transformer. Since LIC-assisted Indices Map Prediction is a form of conditional generation and high-level global priors are already fused into the masked locations via the $\textbf{v}{joint_{P}}$ embedding, we find it beneficial to further constrain the prediction output to more closely align with the ground truth. This is achieved by integrating lower-level fidelity features from the enhanced LIC output $\hat{\textbf{x}}_{lic}$ into the transformer decoder. Specifically, inspired by the “Prediction Assistance” design in HybridFlow [lu2024hybridflow], we use a forked VQ Encoder $\textbf{E}_{P}$, with the output channel $\textbf{C}_{vq}$ modified to match the hidden dimension $h_{dim}$ of the transformer, to extract a lower-level feature map $\textbf{fm}_{P}\in\mathbb{R}^{\frac{H}{n}\times\frac{W}{n}\times h_{dim}}$ from $\hat{\textbf{x}}_{lic}$. The feature map $\textbf{fm}_{P}$ is then flattened to the shape $\frac{HW}{n^{2}}\times h_{dim}$ and fed into the cross-attention modules of the transformer decoder, serving as the keykey and valuevalue.
[…]
3.2.2 LIC-assisted conditioned pixel decoding
@@
Figure 4: VQ Correction Module for dual-stream merging.
It is worth mentioning that our method has the identical bit consumption as the dual-stream HybridFlow [lu2024hybridflow], since both the ground-truth DRV $\textbf{v}_{gt}$ for LIC stream enhancement and latent $\textbf{v}_{join_{P}}$ for improved indices prediction are reconstructed in the decoder. In other words, by learning diffusion priors, we enhance dual-stream performance without additional transmission overhead.
3.3 Training Pipeline
@@
3.3.1 Basic flow pre-training
The LIC stream uses the pre-trained LIC encoder from MLIC. The VQ-Encoder $\textbf{E}_{vq}$ and the learned visual codebook C use the pre-trained VQGAN model [esser2021taming]. These pre-trained components are designed to either pursue high-fidelity reconstruction or high-quality reconstruction, ensuring the baseline performance of our dual-stream system.
3.3.2 DRV-based enhancement module training
@@
3.3.3 DRV-based transformer predictor training
@@
3.3.4 VQ correction module training
@@
3.3.5 DRV-diffusion module training
@@
Figure 6: Qualitative Comparison of our method to the baselines. "1_4" mask strategy ($75\%$ mask ratio on $\textbf{d}_{vq}$) is utilized to maintain around 0.035 bpp within the similar range of the compared baselines. Zoom in for better visualization.
4 Experiments
Datasets. Our HDCompression model is trained on ImageNet-1k [imagenet], with 1000 categories. In alignment with HybridFlow [lu2024hybridflow], performance is evaluated over three benchmarks: Kodak [kodak], CLIC 2020 test set [CLIC2020] and Tecnick dataset [tecknick].
Model configurations. Training images are cropped into $256\!\times\!256$ patches. To achieve approximately 0.025 bpp using pre-trained MLIC models [mlic2023] as the Base LIC, which typically provide a minimum of 0.1 bpp, input patches for the LIC stream are further downsized to $128\times 128$ via bilinear interpolation. Both streams have the same downsampling factor $m=n=16$. The lightweight diffusion modules use a 4-step DDPM scheduler. We unify the loss weights as follows: $w_{1}=1.2$ for $L_{1}$ pixel loss, $w_{2}=0.8$ for AlexNet-based LPIPS perceptual loss, and $w_{3}=0.12$ for UNet-based pixel-wise discriminator GAN loss.
Compared baselines. We compare with several representative image compression methods: 1) Traditional hand-crafted VVC & BPG; 2) MLIC [mlic2023], a single-streamed conventional LIC method; 3) VQGAN [esser2021taming] & Fine-tuned VQGAN [VQPeking], single-streamed VQ-codebook-based methods; and 4) HybridFlow [lu2024hybridflow], a dual-stream framework that straightforwardly combines pre-trained LIC with VQ-codebook-based stream. To obtain ultra-low bitrates, we set QP ranging in $[45,51]$ for VVC & BPG, and downsize the input image (1/2 width and height) for MLIC.
Evaluation metrics. We evaluate commonly used PSNR and LPIPS [LPIPS]. PSNR measures pixel-level distortion and LPIPS assesses the visual quality.
4.1 Quantitative Results
HDCompression has the same bpp as HybridFlow, which operates over $[0.025,0.065]$ bpp range, spanning from fully masked indices map (lowest quality) to unmasked indices map (highest quality). As shown in Fig. 5, traditional handcrafted VVC & BPG and conventional MLIC outperform codebook-based methods over PSNR, due to their learning target of minimizing pixel-level distortions. However, codebook-based methods perform significantly better for perceptual LPIPS. The distortion-driven focus gives artificially inflated PSNR, which omits image details and prefers overly smoothed regions. In contrast, single-stream codebook-based methods emphasize LPIPS, neglecting fidelity to the original image, resulting in >>4 dB PSNR drop compared with traditional methods at the same bpp.
@@ 4.1 Quantitative Results
4.2 Qualitative results
@@
4.3 Ablation Study on Hybrid-Diffusion Modules
@@
4.3.1 DRV-based enhancement for the LIC stream
We compare our LIC stream output ${\hat{\textbf{x}}}_{lic}$ of incorporating the enhancement module against the original output $^{\hat{\textbf{x}}}$ of the pre-trained MLIC. As illustrated in Fig. 7 (a), the enhancement module improves the quality of the original LIC output $^{\hat{\textbf{x}}}$, particularly by reducing blurs. The enhancement module significantly enhances LPIPS of $^{\hat{\textbf{x}}}$ while maintaining almost the same PSNR across various datasets in Fig. 7 (b).
4.3.2 DRV-based transformer for mask prediction
We compare the logit-wise token prediction loss (Eq. 10) of the DRV-based transformer mask predictor against the naive transformer mask predictor in HybridFlow without using DRV. As shown in Fig. 7 (b) where the indices map is masked by "1_4" masking schedule (75% masking ratio), the prediction loss is dropped by 18.5% on average by using DRV in the transformer encoder, leading to 15% accuracy improvement in token prediction on average. Thus, more ground-truth indices are recovered to provide more specific details to the generative stream that might be neglected by the LIC stream.
4.3.3 VQ correction for dual-stream merging